OpenStudy (anonymous):

solve for z: y^2 + 3yz - 8z - 4x = 0

myininaya (myininaya):

\[y^2+z(3y-8)-4x=0\] \[z(3y-8)=4x-y^2\]

myininaya (myininaya):

\[z=\frac{4x-y^2}{3y-8}\]

myininaya (myininaya):

\[y \neq \frac{8}{3}\]

OpenStudy (anonymous):

thanks!

myininaya (myininaya):

np